OpenStudy (anonymous):

Trey's gas tank is 1/2 full. After he buys 5 gallons of gas, it is 5/6 full. How many gallons can Trey's tank hold?

OpenStudy (welshfella):

5/6 - 1/2 = 5/6 - 3/6 = 2/6 = 1/3 of the tank 1/3 of the tanks capacity = 5 gallons so how many gallons does the whole tank hold?

hero (hero):

let × = gallons of full tank. Then we can solve the following equation for x: \[(1/2)x +5= (5/6)x\]